OpenStudy (anonymous):

I have to write an equation in standard form for the ellipse with center (0,0) co-vertex (4,0) and focus (0,3) this really confuses me,

OpenStudy (tkhunny):

Center at (0,0) -- This simplifies things a lot. Focus at (0,3) -- Aha!!!! There must be another focus at (0,-3) ==> c = 3 Co-Vertex at (4,0) -- Aha!!!! There must be another co-vertex at (-4,0) ==> b = 4 There is a relationship between 'a', 'b' and 'c' for an ellipse. Do you know what it is?

OpenStudy (anonymous):

isnt it c^2=a^2+b^2?

OpenStudy (tkhunny):

It is important to be confident on this. Please compare it to the formula for an hyperbola. Make sure you know which is which. For an ellipse, a > c!

OpenStudy (anonymous):

so it would be y2/16+x^2/25=1?

OpenStudy (tkhunny):

I believe you have the 16 and 25 switched. Look carefully at the vertices. Where are they? Which is a greater distance from the origin?